5. StartCo is an early stage company whose financial plans call for the company to be sold in 5 years, at a valuation of $10 million. You are considering an investment of $100,000 in StartCo;

you like the company but feel that it is a fairly risky venture, so you want a 40% annual return on your investment.

a. What is the present value of StartCo, using your required rate of return as the discount rate?

b. What % of the company would you expect to receive for your $100,000?

c. If there are 750,000 shares of StartCo already outstanding, out of 10 million shares authorized, how many shares will there be after you invest your $100,000?

d. How many shares will you be purchasing, and at what price?
Answer #1

A)Present value of the company= expected value after five years/(1+ discount rate)time

=$10m/(1.40)5

=$1.8593million

B) % of the company would we received by investing $100000.

={investment / (Present value of the company +investment) }*100

={$0.1million/(1.8593+0.1)million}*100=5.1039%

C)total no of the shares after investment= 750000/(100-5.1039)*100=790338(rounded off)

D) Price of per share =1.8593million/.75million=2.4791

No of the share to be issued=100000/2.4791=40338(rounded off)